# Beyond cognacy

### v1

@@ Abstract
Beyond cognacy
Gerhard Jäger University of Tübingen, Germany gerhard.jaeger@uni-tuebingen.de June 17, 2025thanks: to appear in Proceedings of The 7th Workshop on Research in Computational Linguistic Typology and Multilingual NLP
@@ 1 Introduction
Originally developed in computational biology, quantitative methods for phylogenetic reconstruction using likelihood-based inference frameworks have now gained widespread acceptance in comparative linguistics. This is evident from the growing number of computational phylogenies proposed for some of the world’s largest language families, including Dravidian (Kolipakam et al., 2018), Sino-Tibetan (Sagart et al., 2019), and Indo-European (Heggarty et al., 2023). Moreover, fully automated approaches — where even cognate identification is performed algorithmically — have demonstrated a surprising degree of robustness (Rama et al., 2018). In contrast to the pre-computational era of historical linguistics, where such detailed reconstructions were rare, the generation of fully resolved phylogenies with branch lengths and, in some cases, estimated divergence dates has now become a standard practice in studies of language evolution.
@@ 2.1 Materials
Word lists were obtained from Lexibank [^1] https://github.com/lexibank; List et al. 2023). These datasets contain lexical entries, including the language they belong to, their meaning, form in IPA transcription, and often a manual cognate annotation. The datasets are curated by the Lexibank community and are available in a standardized format, which makes them suitable for computational analyses.
@@ 2.2.3 Multiple sequence alignment (MSA)
c -0.86
j -1.17
L -1.54
1 -2.94
I -8.06
h -9.37
7 -9.60
i -10.14
y -10.24
T -10.33
Table 2: The ten sound classes with the highest probability of being matched with a gap in the trained pHMM, along with their log-probabilities.
@@
babae 98.26
babau 96.31
bIba 95.73
bawa 87.55
zaba 85.51
raba 74.58
maba 73.52
eaba 73.50
xaba 71.78
naba 70.94
@@
In a fifth step, the MSA was converted to a binary matrix. Two binarization methods were used simultaneously. For a given column in an MSA, a character was created for the presence of a sound class. For column 4 in Table 4, e.g., this character has value $1$ for Nanai, Orok and Ulch, and 00 for the other languages. Additionally, for each sound class type in a column, a different character was created. In the example, there are two such characters, one for k and one for q. The first has value $1$ for Nanai and Orok and 00 otherwise, while the second has value $1$ for Ulch and 00 otherwise. Languages for which the data do not contain a reflex for a given concept are coded as missing for all relevant characters. If a language has multiple reflexes for a given concept, the maximum value is chosen.
Applying this workflow to all concepts and concatenating the resulting matrices yields the final character matrix 928 rows and 46,409 columns.
@@ 2.2.4 Phylogenetic inference
We performed phylogenetic inference using raxml-ng (Kozlov et al., 2019), which implements maximum-likelihood estimation. The GTR+G model (generalized time-reversible model with gamma-distributed rates) was used for all analyses. This means that gain rates and loss rates can be different, and that the mutation rates of the different characters can differ but are drawn from the same gamma distribution. The parameters of this distribution are estimated from the data.
@@ 3 Results
Family Metric cc pmi msa
Afro-Asiatic GQD 0.455 0.195 0.211
PhyDiff 0.420 0.260 0.100
AIC 12.635 8.841 -21.476
@@
PhyDiff 0.250 0.110 0.080
AIC 0.868 -0.845 -0.023
Atlantic-Congo GQD 0.150 0.220 0.235
PhyDiff 0.680 0.670 0.530
AIC -34.076 15.572 18.504
Austroasiatic GQD 0.052 0.093 0.093
PhyDiff 0.280 0.310 0.280
AIC 11.338 -16.115 4.777
Austronesian GQD 0.161 0.200 0.194
PhyDiff 0.670 0.600 0.470
AIC -31.419 65.299 -33.880
Chibchan GQD 0.118 0.332 0.310
PhyDiff 0.350 0.360 0.110
AIC -0.168 -3.895 4.063
Dravidian GQD 0.312 0.242 0.242
PhyDiff 0.390 0.170 0.100
AIC 1.196 -0.862 -0.334
Indo-European GQD 0.031 0.014 0.005
PhyDiff 0.320 0.210 0.140
AIC 1.065 -19.079 18.015
Pama-Nyungan GQD 0.178 0.359 0.299
PhyDiff 0.540 0.420 0.360
AIC -17.375 13.601 3.774
Sino-Tibetan GQD 0.230 0.318 0.279
PhyDiff 0.590 0.300 0.280
AIC 32.455 -13.239 -19.216
@@
PhyDiff 0.270 0.030 0.010
AIC -1.364 0.758 0.607
Tupian GQD 0.353 0.294 0.266
PhyDiff 0.390 0.200 0.160
AIC -0.187 -0.280 0.467
@@
PhyDiff 0.350 0.230 0.170
AIC -1.266 0.647 0.618
Uto-Aztecan GQD 0.126 0.084 0.083
PhyDiff 0.120 0.050 0.050
AIC 2.098 -2.485 0.388
@@ Acknowledgments
This research was supported by the DFG Centre for Advanced Studies in the Humanities Words, Bones, Genes, Tools (DFG-KFG 2237) and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement 834050).

### v2

# Gerhard Jäger

@@ Abstract
Beyond cognacythanks: This is an updated version of the paper published in the Proceedings of SIGTYP 2025, correcting a sign error in the pHMM training code and two bugs in the T-Coffee MSA construction. The qualitative conclusions are unchanged.
Gerhard Jäger University of Tübingen, Germany gerhard.jaeger@uni-tuebingen.de March 27, 2026thanks: Published in Proceedings of The 7th Workshop on Research in Computational Linguistic Typology and Multilingual NLP
@@ 1 Introduction
Originally developed in computational biology, quantitative methods for phylogenetic reconstruction using likelihood-based inference frameworks have now gained widespread acceptance in comparative linguistics. This is evident from the growing number of computational phylogenies proposed for some of the world's largest language families, including Dravidian (Kolipakam et al., 2018), Sino-Tibetan (Sagart et al., 2019), and Indo-European (Heggarty et al., 2023). Moreover, fully automated approaches — where even cognate identification is performed algorithmically — have demonstrated a surprising degree of robustness (Rama et al., 2018). In contrast to the pre-computational era of historical linguistics, where such detailed reconstructions were rare, the generation of fully resolved phylogenies with branch lengths and, in some cases, estimated divergence dates has now become a standard practice in studies of language evolution.
@@ 2.1 Materials
Word lists were obtained from Lexibank [^1] https://github.com/lexibank; List et al. 2023. These datasets contain lexical entries, including the language they belong to, their meaning, form in IPA transcription, and often a manual cognate annotation. The datasets are curated by the Lexibank community and are available in a standardized format, which makes them suitable for computational analyses.
@@ 2.2.3 Multiple sequence alignment (MSA)
2 -0.19
1 -2.09
3 -3.06
q -6.12
E -6.47
5 -7.05
x -7.73
i -9.38
7 -9.79
G -10.64
Table 2: The ten sound classes with the highest probability of being matched with a gap in the trained pHMM, along with their log-probabilities. In List's version of ASJP, numerals 1–6 encode tone contours (level, rising, falling, etc.) and 7 encodes the glottal stop ().
@@
bIba -49.37
babae -49.44
babau -52.13
eaba -62.23
bawa -62.79
zaba -66.39
raba -76.10
xaba -77.01
maba -79.03
naba -90.25
@@
In a fifth step, the MSA was converted to a binary matrix. Two binarization methods were used simultaneously. For a given column in an MSA, a character was created for the presence of a sound class. For column 4 in Table 4, e.g., this character has value $1$ for Nanai, Orok and Ulch, and 0 for the other languages. Additionally, for each sound class type in a column, a different character was created. In the example, there are two such characters, one for k and one for q. The first has value $1$ for Nanai and Orok and 0 otherwise, while the second has value $1$ for Ulch and 0 otherwise. Languages for which the data do not contain a reflex for a given concept are coded as missing for all relevant characters. If a language has multiple reflexes for a given concept, the maximum value is chosen.
Applying this workflow to all concepts and concatenating the resulting matrices yields the final character matrix 928 rows and 54,355 columns.
@@ 2.2.4 Phylogenetic inference
We performed phylogenetic inference using raxml-ng (Kozlov et al., 2019), which implements maximum-likelihood estimation. The BIN+G model (binary substitution model with gamma-distributed rates) was used for all analyses. This means that gain rates and loss rates can be different, and that the mutation rates of the different characters can differ but are drawn from the same gamma distribution. The parameters of this distribution are estimated from the data.
@@ 3 Results
Family Metric $c$ pmi msa
Afro-Asiatic GQD 0.455 0.195 0.212
PhyDiff 0.420 0.260 0.200
AIC 12.635 8.841 -29.519
@@
PhyDiff 0.250 0.110 0.120
AIC 0.868 -0.845 1.088
Atlantic-Congo GQD 0.150 0.220 0.242
PhyDiff 0.680 0.670 0.300
AIC -34.076 15.572 5.987
Austroasiatic GQD 0.052 0.093 0.095
PhyDiff 0.280 0.310 0.270
AIC 11.338 -16.115 -3.022
Austronesian GQD 0.161 0.200 0.160
PhyDiff 0.670 0.600 0.370
AIC -31.419 65.299 -28.045
Chibchan GQD 0.118 0.332 0.360
PhyDiff 0.350 0.360 0.290
AIC -0.168 -3.895 3.884
Dravidian GQD 0.312 0.242 0.261
PhyDiff 0.390 0.170 0.200
AIC 1.196 -0.862 -0.343
Indo-European GQD 0.031 0.014 0.016
PhyDiff 0.320 0.210 0.260
AIC 1.065 -19.079 8.237
Pama-Nyungan GQD 0.178 0.359 0.251
PhyDiff 0.540 0.420 0.450
AIC -17.375 13.601 1.963
Sino-Tibetan GQD 0.230 0.318 0.245
PhyDiff 0.590 0.300 0.260
AIC 32.455 -13.239 -23.957
@@
PhyDiff 0.270 0.030 0.040
AIC -1.364 0.758 0.564
Tupian GQD 0.353 0.294 0.255
PhyDiff 0.390 0.200 0.180
AIC -0.187 -0.280 -0.001
@@
PhyDiff 0.350 0.230 0.120
AIC -1.266 0.647 0.612
Uto-Aztecan GQD 0.126 0.084 0.076
PhyDiff 0.120 0.050 0.110
AIC 2.098 -2.485 0.309
@@ Acknowledgments
This research was supported by the DFG Centre for Advanced Studies in the Humanities Words, Bones, Genes, Tools (DFG-KFG 2237) and by the European Research Council (ERC) under the European Union's Horizon 2020 research and innovation programme (Grant agreement 834050).